OpenStudy (anonymous):

find the value of sin^-1 0 a.-1 b.0 c.1 d.pi

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you know a number whose sine is zero?

OpenStudy (anonymous):

hey misty!!

OpenStudy (misty1212):

what number (or angle) has sine zero?

OpenStudy (anonymous):

i dont actually

OpenStudy (misty1212):

oh you need a nice unit circle cheat sheet hold one

OpenStudy (misty1212):

look at picture of the unit circle on the last page then look at the points one the unit circle the first coordinate is cosine, the second coordinate is sine

OpenStudy (anonymous):

is it 90 degrees?

OpenStudy (misty1212):

for example you can see that at \(\frac{\pi}{2}\) you have the point \((0,1)\) that means \[\cos(\frac{\pi}{2})=0\] and \[\sin(\frac{\pi}{2})=1\]

OpenStudy (misty1212):

you want sine to be zero, not cosine, so no, it is not 90 degrees you want the second number to be zero

OpenStudy (anonymous):

righttt i see that now

OpenStudy (misty1212):

also from your answer choices you are not working with degrees, you are working with radians

OpenStudy (misty1212):

do you see it now?

OpenStudy (anonymous):

yes i believe i do thanks

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

oh be careful, there are lots of places it is zero

OpenStudy (misty1212):

don't pick \(\pi\) that is not correct, pick \(0\)

OpenStudy (anonymous):

yea i just notice that on the unit circle

OpenStudy (anonymous):

ok imma go wit zero

OpenStudy (misty1212):

for the inverse cosine your answer has to be between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\)

OpenStudy (misty1212):

yes, 0 is between those number, \(\pi\) is not

OpenStudy (anonymous):

its correct to. thanks again misty :)